A ray of light passing FROM water TO air plate at an angle of 30° with the normal. Determine


the


a. angle between the refracted ray and


b. the reflected ray.


(please see the table provided)


(With solution)


1
Expert's answer
2021-05-16T17:59:57-0400


First, find the angle of refraction using Snell's law:


"n_w=\frac{\sin\theta_r}{\sin\theta_i},\\\space\\
\theta_r=\arcsin(n_w\sin\theta_i)=42°."

So, the angle between the refracted ray (42° to the normal in air) and the reflected ray (30° to the normal in water) is


"\alpha=(90°-\theta_r)+(90°-\theta_i)=108°."